[Brahms-l] Shift report 20040315 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Mon Mar 15 2004 - 08:35:21 EST
------------------------------------------------------------
Date: Mon Mar 15 08:35:21 2004, Shift: 00:00-08:00
Supervisor: CH, with: 
Log book#15 pages 88 - 88

Shift Summary:
------------------------------------------------------------
1: Horrible!  (or that is, not really the machine, but ...)
   The ZDCs was out of sync with the RHIC clock for the 4
   first hours.  Apparently, the spill was shifted by a 
   bunch, which lead somehow to a change in the clock.

   Had to call Flemming at around 1AM - sorry about that    
   Flemming. 

   From then on, approx 3.5 H of stable beam. 

   Toward the end, there was a Quench Link Interlock.

2: 11032   67k tr 2, 86k tr 3
   11028   10k tr 2, 15k tr 3
   11029   99k tr 2, 74k tr 3

3: The RunServer keeps dying - I think it's due to the pii's 
   cannot contact the NFS server at `opus'.  Some sort of 
   stale handles or something.  Perhaps `opus' needs a once 
   over?  It started when we had the lab-wide network 
   problem.    
   
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Mar 15 08:35:29 2004

This archive was generated by hypermail 2.1.8 : Mon Mar 15 2004 - 08:35:50 EST